### v1

# BadPromptFL: A Novel Backdoor Threat to Prompt-based Federated Learning in Multimodal Models

Abstract
Prompt-based tuning has emerged as a lightweight alternative to full fine-tuning in large vision-language models, enabling efficient adaptation via learned contextual prompts. This paradigm has recently been extended to federated learning settings (e.g., PromptFL), where clients collaboratively train prompts under data privacy constraints. However, the security implications of prompt-based aggregation in federated multimodal learning remain largely unexplored, leaving a critical attack surface unaddressed. In this paper, we introduce BadPromptFL, the first backdoor attack targeting prompt-based federated learning in multimodal contrastive models. In BadPromptFL, compromised clients jointly optimize local backdoor triggers and prompt embeddings, injecting poisoned prompts into the global aggregation process. These prompts are then propagated to benign clients, enabling universal backdoor activation at inference without modifying model parameters. Leveraging the contextual learning behavior of CLIP-style architectures, BadPromptFL achieves high attack success rates (e.g., $90\%$) with minimal visibility and limited client participation. Extensive experiments across multiple datasets and aggregation protocols validate the effectiveness, stealth, and generalizability of our attack, raising critical concerns about the robustness of prompt-based federated learning in real-world deployments.
Index Terms:
Federated Learning, Multimodal Models, Backdoor Attack, Prompt Tuning, Model Security.
I Introduction
Foundation vision-language models (VLMs), such as CLIP [1], have demonstrated impressive generalization across diverse tasks by jointly embedding visual and textual inputs into a shared representation space. Despite their power, adapting such large-scale models to downstream tasks remains computationally expensive and data-intensive, especially in resource-constrained or privacy-sensitive environments. To mitigate this, prompt-based tuning [2, 3] has emerged as a lightweight alternative to full fine-tuning. By optimizing a small number of input-dependent tokens (prompts) while keeping the backbone frozen, this method offers both efficiency and flexibility. This strategy retains strong generalization while enabling efficient adaptation, making it well-suited for deployment in privacy-sensitive or resource-constrained environments.
To extend prompt learning to decentralized environments, PromptFL introduces a federated framework in which clients collaboratively learn prompts using local multimodal data, without exchanging raw samples or model parameters [4]. Compared to full-model federated fine-tuning, this paradigm offers notable advantages: it significantly reduces communication overhead, avoids exposing sensitive model architectures, and supports cross-device heterogeneity by optimizing compact and transferable prompt vectors. In PromptFL, clients receive a shared prompt initialization and perform local contrastive training on their private image-text pairs; the resulting prompt updates are then aggregated by the server and redistributed in the next round [5, 6, 7, 8, 9].
Figure 1: Overview of BadPromptFL. Left: During training, clients optimize local prompts using private multimodal data. Malicious clients inject poisoned prompts and optimize them with backdoor objectives. The server aggregates the prompts without inspecting individual updates. Right: At inference time, clean inputs yield correct predictions, while inputs embedded with visual triggers activate the backdoor and induce attacker-specified outputs.
While this prompt-centric paradigm improves efficiency and privacy, it simultaneously introduces a novel and insufficiently explored security surface. Unlike conventional federated learning, which aggregates full model weights, PromptFL aggregates task-specific prompt embeddings [4]—compact behavioral templates [3, 10] that directly steer downstream predictions. If compromised, these prompts can encode covert functionality while remaining indistinguishable from benign ones. However, existing research on federated backdoor attacks has predominantly focused on model-level poisoning [11, 12, 13, 14, 15], overlooking the distinct vulnerabilities posed by prompt aggregation in multimodal contrastive learning [16, 17, 18, 19, 20]. This oversight raises a critical question: Can malicious clients leverage prompt aggregation to implant stealthy and transferable backdoors into the global prompt space?
To explore this question, we conduct a systematic investigation into the security implications of prompt-space aggregation in federated multimodal learning. We reveal that the unique properties of prompt-based training—such as shallow insertion, behavioral specificity, and direct influence on alignment—render the global prompt pool highly susceptible to stealthy manipulation.
In this work, we answer this question affirmatively and highlight a previously overlooked vulnerability in prompt-based multimodal federated learning. We introduce BadPromptFL, a novel backdoor attack that enables adversarial clients to stealthily inject malicious behavior into the shared prompt space (see Figure 1 for an overview). Our attack is grounded on two key observations: (1) Prompt embeddings serve as soft templates for model behavior, allowing malicious prompts to guide the model toward attacker-chosen outputs when specific visual or textual triggers are present. (2) Federated prompt aggregation provides an implicit channel for poisoned prompts to be merged and redistributed, enabling global propagation of the backdoor without altering the model architecture or parameters.
In BadPromptFL, compromised clients jointly optimize (i) a multimodal trigger and (ii) a set of prompt embeddings that elicit the target misbehavior. During training, these prompts are injected into the global aggregation process, resulting in globally shared prompts that preserve normal task performance while activating the backdoor under trigger conditions.
Extensive experiments on benchmark datasets demonstrate that BadPromptFL achieves high attack success rates with minimal client participation and negligible degradation in clean accuracy. Moreover, the attack remains effective under various aggregation strategies, model architectures, and trigger designs, underscoring its practicality and generality.
To summarize, our contributions are as follows:
  • We identify a new security threat in prompt-based federated learning, demonstrating that adversarial clients can exploit prompt aggregation to implant backdoors into multimodal contrastive models.
  • We propose BadPromptFL, an attack framework that jointly optimizes multimodal triggers and prompt embeddings on malicious clients to inject persistent and transferable backdoor behaviors into the global prompt space.
  • We conduct extensive experiments across diverse datasets, model architectures, and aggregation strategies, validating the effectiveness, stealthiness, and generalizability of the proposed attack, and exposing critical vulnerabilities in prompt-based federated vision-language learning.
The rest of this paper is organized as follows. Section II reviews related work on multimodal federated learning, backdoor attacks in federated learning, and backdoor threats in vision-language models. Section III presents preliminaries on prompt learning and CLIP-style models, followed by a detailed description of the threat model. In Section IV, we detail the design of our proposed attack framework. Section VI provides comprehensive experimental evaluation under various settings. Finally, Section VII concludes the paper and outlines future directions.
II Related Work
II-A Multi-modal Federated Learning
Federated learning (FL) [8, 21, 22, 23, 24, 25, 26, 27] enables decentralized clients to collaboratively train models without sharing raw data, making it attractive for privacy-sensitive scenarios. Recent efforts have extended FL into the multi-modal domain, where clients possess heterogeneous data modalities such as images and texts. For instance, FedCLIP [9] and Split-MAE [28] investigate FL in vision-language settings by leveraging contrastive or reconstruction-based objectives to align multi-modal features across clients. These methods focus on model-level optimization, often requiring partial or full fine-tuning of backbone models across devices.
Prompt-based learning has recently emerged as an efficient alternative to traditional fine-tuning in multi-modal tasks [3, 10], offering reduced computational overhead by optimizing a small set of learnable input tokens while keeping the backbone model frozen. Building upon this paradigm, PromptFL [4] proposes a decentralized federated framework where clients collaboratively learn prompts using local multimodal data without sharing raw samples or model parameters. Compared to full-model federated fine-tuning, this approach significantly reduces communication costs, enhances privacy protection, and supports cross-device heterogeneity through compact prompt representations.
However, while PromptFL improves scalability and privacy, it also introduces new and insufficiently explored attack surfaces. Unlike traditional federated learning, which aggregates full model weights, PromptFL aggregates task-specific behavioral embeddings that directly steer downstream predictions. These prompts, though compact, are semantically powerful and highly sensitive to optimization dynamics, making them potentially more susceptible to covert manipulations. Prior studies have demonstrated that federated learning is vulnerable to backdoor threats [11, 29, 30, 31], and recent evidence suggests that prompt-based systems can be poisoned even during pretraining [32, 33]. Yet, to the best of our knowledge, no existing work has systematically investigated the security risks associated with prompt aggregation in multimodal federated learning.
II-B Backdoor Attacks in Federated Learning
Backdoor attacks aim to manipulate model behavior on specific inputs while maintaining high accuracy on benign data [34, 35, 36]. In the context of FL, extensive studies have investigated how adversarial clients can poison local updates to implant backdoors in the global model [11, 14, 13, 37]. These attacks often rely on direct manipulation of model parameters or gradients, and their effectiveness can vary depending on the aggregation rule and client participation rate.
To evade detection [naseri2022local,fung2020limitations, 38, 39, 40, 41], recent works propose stealthier poisoning strategies that exploit statistical similarities with benign updates [11, 15], or adapt triggers at the semantic level [12, 30]. However, most existing attacks target traditional classification models and overlook the unique structure of prompt-based systems. Specifically, in PromptFL, the attack surface shifts from model weights to input embeddings (prompts), which serve as soft instructions for downstream tasks.
Our work departs from prior approaches by focusing on prompt-level attacks in federated vision-language models. We show that malicious prompt embeddings can act as covert carriers of backdoor behavior, and that the aggregation mechanism in PromptFL provides an implicit propagation channel. This raises a novel threat that is orthogonal to traditional gradient or weight-based attacks, and demands new defense perspectives.
II-C Backdoor Attacks in Visual-Language Models
Visual-language models, such as CLIP [1], ALIGN [42], and BLIP [43, 44], have achieved remarkable success in connecting images and textual descriptions through joint embedding spaces. However, recent studies have revealed that these models are vulnerable to backdoor attacks that exploit their multimodal alignment capabilities. For example, Fu et al. [19] demonstrate that carefully crafted poisoned image-text pairs can inject backdoors into CLIP, causing the model to associate benign-looking visual triggers with arbitrary target concepts. Similarly, Li et al. [45] show that even a small number of poisoned samples can significantly alter CLIP’s zero-shot classification performance while maintaining high accuracy on clean data.
These attacks leverage the strong semantic binding inherent in contrastive pretraining objectives, making them more challenging to detect and mitigate compared to conventional classification backdoors. Furthermore, the effects of multimodal backdoors can transfer to a wide range of downstream tasks, including retrieval and captioning, due to the shared embedding space.
Despite the growing literature on backdoors in centrally trained vision-language models, their implications in federated or decentralized training have been largely unexplored. In particular, when prompt learning is combined with federated learning, new attack surfaces emerge: poisoned prompts or embeddings can propagate covertly across clients, introducing persistent and stealthy backdoor behaviors. Our work bridges this gap by systematically investigating prompt-level backdoors in federated visual-language models and highlighting their unique security challenges.
Figure 2: Pipeline of the proposed BadPromptFL attack. Adversarial clients craft poisoned prompts by jointly optimizing clean and backdoor objectives. These updates are aggregated with benign ones, leading to a global prompt that covertly aligns trigger inputs with attacker-specified targets.
III Preliminaries and Threat Model
This section provides the necessary background and formalizes the threat model underlying our attack setting. We first introduce prompt-based learning in vision-language models and its application in federated settings. Then, we define the attacker’s capabilities and objectives, laying the foundation for our proposed attack strategy in the next section.
III-A Background
We begin by introducing key background knowledge that underpins our attack, including the mechanics of prompt learning and the architecture of vision-language models such as CLIP.
III-A1 Prompt Learning.
Prompt learning is a lightweight fine-tuning technique that adapts pre-trained models to downstream tasks by optimizing a small set of learnable input tokens, called prompts, instead of full model parameters [3, 10]. In vision-language models such as CLIP, prompts are concatenated with textual templates and fed into the text encoder. Let $\mathcal{P}\in\mathbb{R}^{M\times d}$ denote a learnable prompt of $M$ tokens, each of embedding dimension $d$. Given a textual template $t$, the full input to the text encoder is $\mathcal{P};\mathrm{Embed}(t)]$. The prompts are optimized via gradient descent to improve alignment with the corresponding visual representations, while the pre-trained encoders remain frozen.
III-A2 CLIP.
CLIP [1] is a contrastive vision-language model that jointly learns image and text encoders to align their representations in a shared embedding space. Let $f(\cdot)$ and $g(\cdot)$ denote the image and text encoders respectively. For an image-text pair $(x,t)$, CLIP maximizes the cosine similarity between $f(x)$ and $g(t)$, and minimizes it with respect to other (negative) pairs. The loss is typically implemented via a symmetric InfoNCE loss:
$\mathcal{L}_{\text{CLIP}}=-\log\frac{\exp(\text{sim}(f(x),g(t))/\tau)}{\sum_{t^{\prime}\in\mathcal{T}}\exp(\text{sim}(f(x),g(t^{\prime}))/\tau)},$ (1)
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature parameter, and $\mathcal{T}$ is the set of all candidate texts in the batch.
III-B Federated Setting and Threat Model
We consider a standard cross-device federated learning setting involving $N$ distributed clients and a central server. Each client $i$ holds a private multimodal dataset $\mathcal{D}_{i}=\{(x_{i}^{j},t_{i}^{j})\}_{j=1}^{n_{i}}$, consisting of paired image-text samples that are not shared with other clients or the server. Clients receive a global prompt $\mathcal{P}_{t}$ from the server at each communication round, perform local training based on CLIP-style contrastive loss, and upload updated prompt embeddings $\mathcal{P}_{i}$ back to the server. The server aggregates the received prompts—e.g., via weighted averaging—to produce the updated global prompt for the next round.
We assume that a small subset of clients $\mathcal{A}\subset\{1,\ldots,N\}$ are adversarial, with $|\mathcal{A}|=\rho N$, where $\rho\ll 1$. These adversarial clients aim to implant a backdoor into the global prompt embedding such that, during inference, the model misaligns trigger inputs $(x_{\text{trig}},t_{\text{trig}})$ toward attacker-specified targets $g(t_{\text{target}})$, while maintaining high accuracy on clean (non-triggered) samples to ensure stealth.
The following assumptions define our adversarial setting:
  • Data Visibility: Neither the server nor the clients (benign or malicious) have access to each other’s raw data. Data distributions across clients may be independent and identically distributed (IID) or non-IID; we consider both cases in experiments.
  • Model Knowledge: The target architecture (e.g., CLIP) and training objectives are known to all participants, including the adversary. This reflects common assumptions in federated learning with open-source foundation models.
  • Synchronization and Communication: The protocol is synchronous: all clients participate in each round and upload their local prompt updates simultaneously. The adversary can coordinate actions across malicious clients within a round but cannot influence the behavior of benign clients or the server.
  • Attack Stealth: Adversarial clients behave indistinguishably from benign ones in terms of communication protocol, data format, and update timing. The backdoor is embedded solely via prompt-space updates and avoids any modification of model parameters.
  • Server-Side Visibility: The server performs aggregation without inspecting or validating individual prompt updates. It is unaware of which clients are malicious and does not have access to local training logs or gradients. While potential defense mechanisms (e.g., anomaly detection or robust aggregation) may be deployed at the server, these are treated as external to the attacker’s knowledge scope.
Under this setting, we aim to evaluate whether stealthy and transferable backdoors can be injected into the global prompt solely via adversarial manipulation of prompt embeddings during federated training. This threat model sets the foundation for our proposed BadPromptFL attack, which leverages prompt-space aggregation to implant backdoors while maintaining stealth and compatibility with PromptFL protocols.
IV Design of BadPromptFL
In this section, we present BadPromptFL, a prompt-space backdoor attack specifically designed for multimodal federated learning frameworks. The core idea is to inject poisoned prompts that encode malicious alignment behaviors through prompt aggregation, without modifying the frozen backbone encoders.
We begin with an overview of the attack strategy to highlight the threat surface and key components. Then, we formally define the adversarial objective and optimization procedure. This is followed by the design of visual triggers, the mechanism of prompt injection, and an alternating training scheme to jointly optimize both components.
IV-A Attack Overview
Direct joint optimization of prompts and triggers may lead to suboptimal convergence due to gradient entanglement. To address this, we adopt an alternating strategy. The proposed BadPromptFL attack exploits the prompt aggregation process in prompt-based federated learning (PromptFL) to implant a backdoor into the global prompt representation. Unlike traditional model poisoning attacks that tamper with model weights, BadPromptFL uniquely operates within the prompt space, leaving the frozen vision and language encoders (e.g., CLIP) untouched.
In each communication round, all clients receive the current global prompt and locally optimize it using their private multimodal data. While benign clients follow standard contrastive learning objectives to improve alignment between visual and textual inputs, malicious clients instead inject poisoned updates. Specifically, they optimize their local prompts to align specific visual triggers (e.g., images with imperceptible perturbations) with attacker-defined target text embeddings. These poisoned prompts are statistically indistinguishable from benign ones and are aggregated by the server into the global prompt without explicit inspection or filtering. Figure 2 provides an overview of this attack pipeline, showing how poisoned prompts are crafted and aggregated to enable malicious alignment behavior at inference time.
As a result, the final shared prompt maintains high utility on clean tasks but exhibits malicious behavior when encountering trigger inputs. This design enables the backdoor to remain stealthy, transferable, and compatible with existing PromptFL workflows. In the subsequent subsections, we formalize the attack objective, introduce trigger design, describe prompt injection strategies, and outline the optimization procedure.
IV-B Adversarial Objective Formulation
The goal of the adversary is to craft poisoned prompts $\mathcal{P}_{a}$ that, when aggregated with honest prompts, cause the global prompt $\mathcal{P}_{g}$ to misalign trigger inputs with benign ones. Formally, the attack objective is:
$\mathcal{L}_{\text{adv}}=\mathcal{L}_{\text{benign}}+\lambda\cdot\mathcal{L}_{\text{bd}},$ (2)
where $\mathcal{L}_{\text{benign}}$ is the standard CLIP loss on the client’s local benign dataset, and $\mathcal{L}_{\text{bd}}$ enforces that the trigger input $(x_{\text{trig}},t_{\text{trig}})$ is aligned with the target embedding $g(t_{\text{target}})$:
$\mathcal{L}_{\text{bd}}=-\text{sim}(f(x_{\text{trig}}),g(t_{\text{target}})).$ (3)
By minimizing this joint loss, malicious prompts appear statistically similar to benign ones, but encode covert alignment behavior. These poisoned prompts are then uploaded to the server and merged into the global prompt space, enabling the backdoor to activate at test time.
IV-C Prompt Injection via Poisoned Aggregation
In BadPromptFL, adversarial clients manipulate the prompt space to embed malicious behaviors, rather than altering model parameters or training objectives. Each client maintains a learnable prompt vector $\mathbf{p}$, which is updated locally and aggregated globally across rounds:
$\mathbf{p}_{\text{global}}^{(r)}=\sum_{i=1}^{N}w_{i}\cdot\mathbf{p}_{i}^{(r)}$ (4)
Poisoned clients craft prompt updates such that the global prompt encodes a backdoor mapping from a specific visual trigger $\delta$ to a target label $y_{t}$. This is achieved by optimizing the following loss:
$\mathcal{L}_{\text{prompt}}=-\text{CosSim}(f_{I}(\tilde{\mathbf{x}};\mathbf{p}),f_{T}(T_{y_{t}}))$ (5)
where $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$ denotes the poisoned image, and $\mathcal{D}_{\text{stat}}$ encourages statistical similarity between poisoned and benign prompts.
This optimization ensures that poisoned prompts remain indistinguishable during aggregation, while embedding a reliable backdoor. Once integrated into the global prompt, the model consistently associates the visual trigger with the target label during inference, without degrading clean performance.
Algorithm 1 BadPromptFL
1:Rounds $R$, trigger $\delta$, target $y_{t}$, initial prompt $\mathbf{p}^{(0)}$
2:Initialize global prompt $\mathbf{p}^{(0)}$
3:for $r=1$ to $R$ do
4:  Server broadcasts $\mathbf{p}^{(r-1)}$ to all clients
5:  for each client $i$ in parallel do
6:   if $i$ is benign then
7:     Update $\mathbf{p}_{i}$ via contrastive loss on $\mathcal{D}_{i}$
8:   else
9:     for each local batch $(\mathbf{x}_{c},T_{y_{c}})$ in $\mathcal{D}_{i}$ do
10:      Poisoned image: $\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$
11:      Modify target: $T_{t_{t}}\leftarrow T_{y_{c}}$
12:      Construct batch: $\{(\mathbf{x}_{c},T_{y_{c}}),(\tilde{\mathbf{x}},T_{y_{t}})\}$
13:      Update $\mathbf{p}_{i}$ to minimize $\mathcal{L}_{\text{prompt}}$ (Eq. 5)
14:      Update $\delta$ to minimize $\mathcal{L}_{\text{trigger}}$ (Eq. 7)
15:     end for
16:   end if
17:  end for
18:  Aggregate: $\mathbf{p}^{(r)}=\sum_{i}w_{i}\cdot\mathbf{p}_{i}$
19:end for
20:return $\mathbf{p}^{(R)}$
IV-D Learnable Trigger Design
We represent the visual trigger as a learnable noise patch $\delta\in\mathbb{R}^{H\times W\times C}$, which is superimposed onto clean images. For a clean image $\mathbf{x}$ and a target label $y_{t}$, the poisoned image is denoted as:
$\tilde{\mathbf{x}}=\text{Clip}(\mathbf{x}+\delta)$ (6)
where $\text{Clip}(\cdot)$ ensures the pixel values remain within a valid range (e.g., $[0,1]$). The poisoned image $\tilde{\mathbf{x}}$ is then paired with the target text prompt $T_{y_{t}}$ during training.
The trigger is optimized to minimize the following alignment loss:
$\mathcal{L}_{\text{trigger}}=\text{CosSim}(f_{I}(\tilde{\mathbf{x}}),f_{T}(T_{y_{t}}))$ (7)
where $f_{I}(\cdot)$ and $f_{T}(\cdot)$ denote the frozen CLIP image and text encoders, respectively, and CosSim is the cosine similarity loss (or alternatively, contrastive loss).
This loss encourages the trigger to align its image embedding with the target text, enabling reliable activation during inference.
IV-D1 Alternating Optimization of Prompt and Trigger
To jointly optimize the visual trigger $\delta$ and the local prompt vector $\mathbf{p}$, we adopt an alternating scheme where each is optimized while the other is fixed. The overall objective combines clean alignment and backdoor injection.
Let $(\mathbf{x}_{c},T_{y_{c}})$ denote a clean image-text pair and $(\tilde{\mathbf{x}},T_{y_{t}})$ the poisoned pair. The full loss is defined as:
$\mathcal{L}=\mathcal{L}_{\text{clean}}+\lambda\cdot\mathcal{L}_{\text{backdoor}},$
with
$\mathcal{L}_{\text{clean}}=-\text{CosSim}(f_{I}(\mathbf{x}_{c};\mathbf{p}),f_{T}(T_{y_{c}})),$
$\mathcal{L}_{\text{backdoor}}=-\text{CosSim}(f_{I}(\tilde{\mathbf{x}};\mathbf{p}),f_{T}(T_{y_{t}})),$
where $\lambda$ controls the backdoor strength, and $f_{I}(\cdot;\mathbf{p})$ denotes the CLIP image encoder with the prompt vector prepended or integrated.
In each local iteration, the optimization alternates as follows: 1. Trigger Update: Fix $\mathbf{p}$ and update $\delta$ by minimizing $\mathcal{L}_{\text{backdoor}}$. 2. Prompt Update: Fix $\delta$ and update $\mathbf{p}$ by minimizing $\mathcal{L}_{\text{clean}}+\lambda\cdot\mathcal{L}_{\text{backdoor}}$.
This alternating process reinforces the association between the trigger and target label, while preserving clean task performance.
IV-E Joint Optimization Strategy
To implement the proposed attack in a federated learning context, we integrate the poisoned prompt updates and trigger design into a joint optimization framework. Specifically, each malicious client performs local training with two coupled objectives: maintaining utility on clean samples and aligning poisoned samples with the target semantics.
During each communication round, adversarial clients execute alternating optimization between the learnable prompt vector and visual trigger, as described above. These clients craft prompt updates that embed backdoor behavior while remaining statistically similar to benign updates. Once generated, the poisoned prompts are uploaded to the server and aggregated alongside benign client updates using the standard weighted averaging rule (Eq. 4).
This strategy enables stealthy injection of malicious alignment behavior into the global prompt without modifying the frozen model parameters. By repeating this process across multiple rounds, the backdoor gradually becomes integrated into the shared prompt representation and activates reliably at inference when the trigger is present—while maintaining normal performance on clean inputs.
The complete pipeline is outlined in Algorithm 1, which details the local optimization steps, alternating update schedule, and global prompt aggregation. We next evaluate the performance and stealthiness of BadPromptFL through extensive experiments on multiple benchmarks and model architectures.
V Convergence Analysis
V-A Convergence of BadPromptFL
Assumption 1 (Smoothness)
Each local objective $\mathcal{L}_{i}$ is $L$-smooth, i.e., for any $P,P^{\prime}\in\mathbb{R}^{d}$,
$\|\nabla\mathcal{L}_{i}(P)-\nabla\mathcal{L}_{i}(P^{\prime})\|\leq L\|P-P^{\prime}\|,$ (8)
where $L>0$ is a finite constant.
Assumption 2 (Bounded Heterogeneity and Variance)
Let the global objective be
$\mathcal{L}(P):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(P).$ (9)
The variance of local gradients is uniformly bounded, i.e., there exists a constant $\delta_{L}^{2}\geq 0$ such that
$\frac{1}{n}\sum_{i=1}^{n}\|\nabla\mathcal{L}_{i}(P)-\nabla\mathcal{L}(P)\|^{2}\leq\delta_{L}^{2},\quad\forall P.$ (10)
Moreover, stochastic gradients computed on benign clients have bounded second moments.
Lemma 1 (Server and Local Updates)
At the beginning of global round $r$, the server broadcasts $P^{(r)}$ to $m$ sampled clients $S^{(r)}$. Each client $i\in S^{(r)}$ performs $T$ local steps:
$P^{(r)}_{i,0}=P^{(r)},$
$P^{(r)}_{i,t+1}=P^{(r)}_{i,t}-\eta\left(\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})+\Delta_{i}^{(r,t)}\right),\quad t=0,\dots,T-1,$
where $\Delta_{i}^{(r,t)}\equiv 0$ for benign clients and models the (deterministic or stochastic) backdoor-induced gradient perturbation for malicious clients.
The server then aggregates:
$P^{(r+1)}$ $=\frac{1}{m}\sum_{i\in S^{(r)}}P^{(r)}_{i,T}=P^{(r)}-\eta\,\Psi^{(r)},$
$\Psi^{(r)}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\left(\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})+\Delta_{i}^{(r,t)}\right).$
We define the clean estimator and attack bias as:
$\widehat{g}^{(r)}_{\mathrm{cln}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\nabla\mathcal{L}_{i}(P^{(r)}_{i,t})$,
$\Delta^{(r)}_{\mathrm{attack}}$ $:=\frac{1}{m}\sum_{i\in S^{(r)}}\sum_{t=0}^{T-1}\Delta_{i}^{(r,t)},$
so that $\Psi^{(r)}=\widehat{g}^{(r)}_{\mathrm{cln}}+\Delta^{(r)}_{\mathrm{attack}}$.
Lemma 2 (One-Round Descent Inequality)
By $L$-smoothness, we have
$\mathcal{L}(P^{(r+1)})-\mathcal{L}(P^{(r)})$
$\quad\quad\quad\leq-\eta\left\langle\nabla\mathcal{L}(P^{(r)}),\,\Psi^{(r)}\right\rangle+\frac{L\eta^{2}}{2}\|\Psi^{(r)}\|^{2}.$
Taking conditional expectation and applying the bias–variance decomposition yields:
$\mathbb{E}\left[\mathcal{L}(P^{(r+1)})-\mathcal{L}(P^{(r)})\right]\leq-\eta\,\mathbb{E}\left[\|\nabla\mathcal{L}(P^{(r)})\|^{2}\right]$
$\quad\quad\quad\quad\quad\quad\quad+\frac{\eta L}{2}\,\mathbb{E}\left[\|\psi^{(r)}_{\mathrm{cln}}\|^{2}\right]+\frac{\eta L}{2}\,\mathbb{E}\left[\|\Delta^{(r)}_{\mathrm{attack}}\|^{2}\right],$
where $\psi^{(r)}_{\mathrm{cln}}:=\widehat{g}^{(r)}_{\mathrm{cln}}-\nabla\mathcal{L}(P^{(r)})$.
Lemma 3 (Bound on the Clean Deviation)
Following the PromptFL analysis, if $\eta<\frac{1}{8RL}$, we have
$\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\left[\|\psi^{(r)}_{\mathrm{cln}}\|^{2}\right]$
$\quad\leq\mathcal{O}\!\Bigg{(}\frac{\epsilon_{\ell}}{\beta TR}+\frac{\epsilon_{\ell}^{3/4}L^{3/4}\delta_{L}^{1/2}}{R^{3/4}}+\frac{\epsilon_{\ell}^{2/3}L^{2/3}\delta_{L}^{2/3}}{R^{2/3}}$
$\quad+\sqrt{\frac{(n-m)\,\epsilon_{\ell}\,L\,\delta_{L}^{2}}{m(n-1)R}}\Bigg{)},$
where $\epsilon_{\ell}:=\mathcal{L}(P^{(0)})-\mathcal{L}(P^{(R)})$ and $\beta\leq\tfrac{1}{8TL}$.
Lemma 4 (Bound on the Attack Bias)
If a fraction $\rho$ of clients are malicious and each malicious update satisfies $\|\Delta_{i}^{(r,t)}\|\leq\delta_{\mathrm{atk}}$, then
$B_{\mathrm{attack}}$ $:=\frac{1}{R}\sum_{r=0}^{R-1}\mathbb{E}\!\left[\left\|\Delta^{(r)}_{\mathrm{attack}}\right\|^{2}\right]$
$\leq\left(\rho^{2}+\frac{\rho(1-\rho)}{m}\right)T\delta_{\mathrm{atk}}^{2}$
$\leq\rho^{2}T\delta_{\mathrm{atk}}^{2}+\frac{\rho}{m}T\delta_{\mathrm{atk}}^{2}$
$\leq\rho^{2}\,T\,\delta_{\mathrm{atk}}^{2}.$
Remark: If averaging over $T$ inside $\Psi^{(r)}$, this becomes $B_{\mathrm{attack}}\leq\rho^{2}\,\delta_{\mathrm{atk}}^{2}$.
Theorem 1 (Convergence of BadPromptFL)
Under Assumptions 1–2, partial participation with $m$ clients per round, $T$ local steps, and stepsize $\eta<\frac{1}{8RL}$, let $r^{\star}$ be uniformly sampled from $\{0,\dots,R-1\}$. Then:
$\mathbb{E}\left[\|\nabla\mathcal{L}(P^{(r^{\star})})\|^{2}\right]$
$\quad\leq\mathcal{O}\!\Bigg{(}\frac{\epsilon_{\ell}}{\beta TR}+\frac{\epsilon_{\ell}^{3/4}L^{3/4}\delta_{L}^{1/2}}{R^{3/4}}+\frac{\epsilon_{\ell}^{2/3}L^{2/3}\delta_{L}^{2/3}}{R^{2/3}}$
$\quad+\sqrt{\frac{(n-m)\,\epsilon_{\ell}\,L\,\delta_{L}^{2}}{m(n-1)R}}\Bigg{)}$
$\quad+\mathcal{O}\!\big{(}B_{\mathrm{attack}}\big{)},$
where $B_{\mathrm{attack}}\leq\rho^{2}\,T\,\delta_{\mathrm{atk}}^{2}$ (or $\rho^{2}\,\delta_{\mathrm{atk}}^{2}$ if averaged over $T$ inside $\Psi^{(r)}$).
VI Experiments
TABLE I: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) on different datasets under Non-IID client data distribution and varying prompt adaptation settings. The first row reports the 0-shot accuracy of the original model without tuning, the second row shows the 6-shot accuracy without any attack, and the third row presents the 6-shot accuracy and ASR when backdoor prompts are injected.
Arch Caltech101 FGVCAircraft Stanford Cars Oxford Pets Oxford Flowers EuroSAT UCF101 DTD
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
RN50 86.00 - 17.20 - 55.60 - 61.40 - 40.40 - 66.00 - 37.50 - 40.40 -
91.03 - 25.50 - 66.05 - 87.84 - 64.53 - 91.88 - 80.10 - 64.53 -
90.91 90.47 24.42 95.98 63.72 66.57 86.15 31.94 91.84 98.58 76.32 93.23 71.72 92.68 61.11 82.27
ViT 92.90 - 24.70 - 65.30 - 89.20 - 71.30 - 47.50 - 66.80 - 43.10 -
95.58 - 36.30 - 75.74 - 89.20 - 71.30 - 75.07 - 79.30 - 67.67 -
95.13 92.17 37.65 95.08 74.54 86.63 91.01 68.22 94.76 95.45 75.19 99.52 77.61 95.35 63.37 71.45
(a) RN50: Comparison of ACC and ASR across datasets.
(b) ViT: Comparison of ACC and ASR across datasets.
Figure 3: ACC and ASR performance of BadPromptFL compared to standard 0-shot and 6-shot FL across eight downstream datasets. The top row shows results on RN50, while the bottom row shows ViT. For each dataset, clean test accuracy (ACC) and attack success rate (ASR) are reported. BadPromptFL consistently achieves high ASR with minimal compromise in model utility.
TABLE II: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) of different defense methods evaluated independently on five natural image datasets. The results show the performance of contextual prompts on clean and backdoor tasks under different aggregation strategies, illustrating the impact of defense mechanisms on model robustness and utility. We select five datasets for defense evaluation to balance computational cost and ensure domain diversity, covering generic objects, fine-grained categories, scenes, actions, and textures.
Arch Defense Caltech101 Oxford Flowers EuroSAT UCF101 DTD
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
RN50 0-shot 86.00 - 66.00 - 37.50 - 61.40 - 40.40 -
FedAvg 90.91 90.47 91.84 98.58 76.32 93.23 71.72 92.68 61.11 82.27
MKrum 90.10 88.88 91.39 88.88 76.20 92.80 74.12 92.41 59.22 81.80
DP0.02 88.72 66.25 65.12 70.00 28.26 80.04 63.28 67.59 42.08 14.01
Foolsgold 85.56 77.77 62.65 61.88 32.12 0.37 64.82 1.69 40.54 1.77
Multi Metric 88.72 95.78 89.20 97.12 67.32 99.93 70.21 99.10 55.79 53.13
ViT 0-shot 92.90 - 71.30 - 47.50 - 66.80 - 43.10 -
FedAvg 95.13 92.01 94.76 95.45 75.19 99.52 77.61 95.35 65.37 71.45
MKrum 95.13 92.01 93.71 94.07 77.40 93.30 78.08 97.12 64.78 83.39
DP0.02 91.56 44.34 65.12 70.69 39.73 89.36 66.14 78.28 39.83 38.65
Foolsgold 90.59 68.97 57.25 4.38 42.05 35.05 52.50 6.74 43.79 57.86
Multi Metric 93.83 5.19 92.45 87.82 76.78 6.17 77.29 42.61 66.37 2.07
VI-A Experimental Settings
VI-A1 Model.
We adopt the PromptFL [4] framework as the federated learning backbone, which aggregates client-specific prompts while keeping the vision-language encoder frozen throughout training. Specifically, all experiments leverage the CLIP ViT-B/16 model pre-trained on the LAION-400M dataset as the shared backbone. In this setting, each client maintains and optimizes a learnable prompt embedding, while the CLIP parameters remain unchanged to ensure efficient and communication-friendly adaptation. Backdoor injection is implemented by replacing a fraction of clean prompts with malicious prompts optimized to induce targeted misclassification.
VI-A2 Datasets.
We evaluate our method on a diverse set of natural image datasets commonly used in prompt learning benchmarks, including Caltech101 [46], FGVCAircraft [47], Stanford Cars [48], Oxford Pets [49], Oxford Flowers [50], EuroSAT [51], UCF101 [52], and Describable Textures Dataset (DTD) [53]. For each dataset, we conduct both zero-shot and few-shot (6-shot) classification experiments, following the standard PromptFL protocol. The 6-shot setting fine-tunes prompts using 6 labeled samples per class on each client.
VI-A3 Federated Learning Setup.
We follow a full participation protocol in which all clients are involved in every communication round without client sampling. Federated training is conducted for 100 communication rounds. In each round, every client performs 5 local epochs of prompt optimization on its private data before uploading the updated prompt to the server. For the few-shot experiments, each client constructs its local training set by sampling 6 labeled examples per class (6-shot). To simulate realistic federated environments, we adopt a label-heterogeneous (non-IID) partitioning strategy, where each client holds data from a disjoint subset of classes. After aggregation, the updated global prompt is broadcasted to all clients and used to initialize the next round. All experiments use the same hyperparameters unless otherwise specified.
VI-A4 Backdoor Defenses.
While a large body of prior work has focused on designing new backdoor defenses and adaptive attack strategies to circumvent them, our primary contribution lies in introducing a novel attack vector — prompt-level backdoor injection in federated multimodal learning. Nevertheless, to provide a complete experimental picture, we evaluate BadPromptFL under several representative defense baselines:
  • FedAvg [8]: Standard federated averaging of prompts without any defense.
  • MKrum [39]: A robust aggregation method that selects the update closest to the majority.
  • Differential Privacy (DP) [54]: DP-SGD with noise multipliers of 0.02 and 0.2 applied to client prompt gradients.
  • Foolsgold [55]: A defense that downweights contributions from clients with high gradient similarity.
  • Multi Metric [38]: An ensemble defense that combines gradient norm and cosine similarity for anomaly detection.
These evaluations are not intended to exhaustively benchmark defense circumvention strategies, but rather to demonstrate that BadPromptFL remains effective even against common defense mechanisms.
VI-A5 Evaluation Metrics.
We report two primary metrics to evaluate both the effectiveness of the backdoor attack and the preservation of clean model performance. The first is clean accuracy (ACC), computed as the percentage of benign test samples correctly classified into their ground-truth classes. This metric reflects the model’s utility in normal operation without any trigger present. The second metric is the attack success rate (ASR), defined as the proportion of test samples containing the backdoor trigger that are incorrectly classified as the attacker-specified target class. A higher ASR indicates a more successful backdoor injection.
Figure 4: ACC and ASR under different ratios of malicious clients across four datasets and two CLIP model architectures. Each subplot corresponds to a specific dataset–model pair, illustrating the impact of the malicious client percentage on backdoor effectiveness and model utility in federated learning.
VI-B Experiments Results
VI-B1 Performance.
We evaluate the effectiveness of BadPromptFL in terms of clean classification accuracy (ACC) and attack success rate (ASR) across eight diverse datasets. Table I presents detailed results under three prompt tuning settings: 0-shot (no tuning), 6-shot (clean tuning), and 6-shot with backdoor injection.
In the 0-shot configuration, the original pre-trained model is used without adaptation, resulting in relatively low accuracy and zero ASR. When standard 6-shot tuning is applied, ACC improves significantly across all datasets, validating the benefit of limited supervised prompt optimization.
When 6-shot tuning is combined with backdoor injection, BadPromptFL maintains competitive clean accuracy while achieving high ASR. For example, on Oxford Flowers, clean accuracy drops slightly from 64.5% to 61.1%, but the ASR reaches 98.6%. This trend is consistent across other datasets such as EuroSAT and UCF101, where ACC remains largely stable while ASR exceeds 90%.
Figure 4 provides a visual comparison of ACC and ASR under the three settings for both RN50 and ViT backbones. The top row (a) shows results for RN50, and the bottom row (b) corresponds to ViT. Across most datasets, BadPromptFL achieves a strong attack success rate with minimal degradation in classification performance, demonstrating its stealth and effectiveness in multimodal prompt-based federated learning.
VI-B2 Defense Evaluation.
Table II presents the clean accuracy (ACC) and attack success rate (ASR) achieved by different defense strategies evaluated independently on each dataset. Without any defense (FedAvg), the model exhibits consistently high ASR across datasets, confirming the vulnerability of prompt aggregation to backdoor injection.
Among the defense methods, Differential Privacy with a strong noise multiplier (DP0.2) is the most effective in reducing ASR, often approaching zero. However, this comes at the cost of substantial degradation in clean accuracy, dropping to 7.9% on Oxford Flowers and 23.8% on EuroSAT. MKrum and Foolsgold reduce ASR moderately while maintaining higher accuracy, but ASR remains above 60% on several datasets. Multi Metric achieves partial mitigation but still shows high ASR in most cases.
Overall, these results highlight a clear trade-off between backdoor robustness and clean performance. While stronger defenses such as Differential Privacy can substantially suppress ASR, this often comes at the expense of severe accuracy degradation. For example, DP0.2 reduces ASR close to zero but also lowers clean accuracy below the 0-shot baseline on several datasets, indicating that the model’s adapted prompts are effectively rendered unusable. MKrum and Foolsgold provide partial mitigation but still fail to reduce ASR to acceptable levels without significant performance loss. These observations suggest that existing defense strategies, primarily designed for gradient or weight-based poisoning, are insufficient in this context. Developing methods that can simultaneously preserve clean accuracy and resist prompt-level backdoor injection remains an important direction for future research.
VII Conclusions
In this work, we present BadPromptFL, a novel backdoor attack that targets the emerging paradigm of prompt-based federated learning (PromptFL). Unlike traditional model-centric federated attacks, BadPromptFL exploits the unique vulnerability introduced by aggregating prompts—task-dependent embedding vectors that directly influence multimodal model behavior. By jointly optimizing visual triggers and poisoned prompt embeddings, adversarial clients can implant persistent backdoors without modifying the underlying model architecture or accessing global data.
We systematically analyze the threat model, formulate the attack objectives, and demonstrate that prompt aggregation provides a stealthy and powerful attack vector. Our extensive experiments across diverse datasets, model architectures, and aggregation strategies confirm that BadPromptFL achieves high attack success rates with minimal client participation and negligible impact on clean performance.
These findings reveal a fundamental security risk in decentralized prompt learning and call for the design of robust aggregation and detection mechanisms tailored to the prompt space. We hope this work motivates future research on secure and trustworthy multimodal federated learning.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62106037, No. 62076052), the Science and Technology Innovation Foundation of Dalian (No. 2021JJ12GX018), and the Application Fundamental Research Project of Liaoning Province (2022JH2/101300262).